OpenStudy (anonymous):

a bacteria colony doubles in 7 hr. how long does it take the colony to triple? use N=N 02t/T where N0 is the initial number of bacteria and T is the time in hours it takes the colony to double. ( round to nearest hundredth as necessary)

OpenStudy (anonymous):

\[2^{\frac{t}{7}}=3\] solve for \(t\)

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

change of base formula gives \[\frac{t}{7}=\frac{\ln(3)}{\ln(2)}\] and so \[t=\frac{7\ln(3)}{\ln(2)}\] and a calculator

OpenStudy (anonymous):

11.09473751

OpenStudy (anonymous):

is it right

OpenStudy (anonymous):

i don't know, let me check

OpenStudy (anonymous):

looks good to me http://www.wolframalpha.com/input/?i=7log%283%29%2Flog%282%29

OpenStudy (anonymous):

ok thanks you very much

OpenStudy (anonymous):

A bit of explanation [N / No] = (2)^(t/7) ln [ ] = (t/7) ln(2) t = 7 ln [ ] / ln(2) ln [3] = 1.10 ln(2) =.693 t = 11.1